OpenStudy (anonymous):

how do I find the maximum of a quadratic function? how do I find the maximum of a function table?

OpenStudy (danjs):

which class

OpenStudy (anonymous):

algebra 1

OpenStudy (triciaal):

the highest y-value find the vertex a is negative f(x) = ax^2 + bx + c vertex form a(x-h)^2 + k vertex (h, k) h = -b/(2 a) and k = f(h)

OpenStudy (anonymous):

oh so if for example the equation is y=-2x^2-3x+4 , the -2 would be the maximum?

OpenStudy (triciaal):